A student drops a rock from a bridge to the
water 12.4 m below.
With what speed does the rock strike
the water? The acceleration of gravity is
9.8 m/s^2

.
Answer in units of m/s
1
Expert's answer
2019-10-22T10:41:49-0400

The speed of the rock, according to the gravitational acceleration is "v = g t". The time it takes to fall from given height "h = 12.4 m" is "t = \sqrt{\frac{2 h}{g}}"(which is found from the equation of motion "y(t) = h - \frac{g t^2}{2}" by equating "y" to zero and solving for t).

Hence, substituting the time into the equation for speed, obtain "v = g t = g \sqrt{\frac{2 h}{g}} = \sqrt{2 g h} \approx 15.6 s".